# Topological phases of extended Su-Schrieffer-Heeger-Hubbard model

### v1

@@ I Introduction
However, in many experiments setups, the interaction between electrons dominates over the hopping amplitude [23, 24, 25], rendering the independent-electron approximation invalid. Topological descriptions of strongly correlated systems are an emerging area of research with significant potential [26, 27, 28]. In interacting systems, the bulk-edge correspondence is modified as the relation between bulk topological property and the many-body, ground-state degeneracy on the edge [29]. Researchers have also developed a range of metrics to evaluate characterize topological edge states, including the Green function method [30], entanglement entropy between topological edges and bulk states [31, 32], and topological phase transition induced by magnetic flux changes [33].
Su-Schrieffer-Heeger-Hubbard (SSHH) model with the nearest neighbour hopping has been extensively studied. However, the one with next-nearest neighbour hopping remains largely unexplored. Here, we discuss the ground-state properties of a one-dimensional SSHH model with next-nearest neighbour hopping. In the absence of Hubbard interaction, this model can be well described by the single-particle theory and exhibits four distinct topological phases ($\mathcal{W}=-1,0,1,2$) without breaking chiral symmetry. According to the bulk-edge correspondence, these topological phases are characterized by the degeneracy of zero-energy edge modes. However, in the presence of onsite Hubbard interaction, we should consider the relation between bulk topological property and the many-body ground-state. We evaluate long-range spin correlation and the second-order Rényi entropy of the ground state, and determine the topological state of the system by measuring the correlation between the boundary and bulk. The presence of Coulomb interaction leads to a Mott gap at half-filling and dimerization at quarter-filling [34]. To mitigate the influence of finite site effects on our results, we employ the constrained-path auxiliary-field quantum Monte Carlo (CP-AFQMC) algorithm. Furthermore, the applicability of constrained-path method extends beyond the half-filling case [32] where there is no sign problem. Our analysis reveals that the system exhibits long-range spin correlation and near-zero Rényi entropy at half-filling, indicative of a topological order state. Furthermore, a long-range antiferromagnetic order emerges at quarter-filling. The application of an external magnetic field disrupts this antiferromagnetic order, leading to a restoration of long-range spin correlation and near-zero Rényi entropy.
@@ III.1 Edge population
Figure 3: Population number of the edge states for $N=8$ sites chain with $J_{11}=J_{33}=0$, and $n$ is the number of electrons. (a) $\sim$ (c) for $U=0$ and (d) $\sim$ (f) for $U=10$. According to Fig. 1(e), the edge states appear at the first boundary ($1_{st}$ and $N_{th}$) sites or the third-boundary ($3_{rd}$ and $(N-2)_{th}$) sites. (a), (d) $n_{edge1}$ is the sum of electron number on the $1_{st}$ and $N_{th}$ sites; (b), (e) $n_{edge2}$ is the sum of electron number on the $3_{rd}$ and $(N-2)_{th}$ sites; (c), (f) $n_{edge3}$ is the sum of electron number on above four sites. In the non-trivial phases without Hubbard $U$, the edge population increases sharply at half-filling, while with Hubbard $U$ the edge population increases sharply at quarter-filling and there-quarter-filling. And in the trivial phase, the edge population increases gradually. There are some wrinkles on the surface in (d) $\sim$ (f), which is due to the finite-size effects.
@@
The difference between nearest neighbour hopping and next-nearest hopping $\Delta_{J_{3}-J_{1}}$ will affect the topological properties of $3_{rd}$ and $(N-2)_{th}$ sites according to Fig. 1(c). When $\Delta_{J_{3}-J_{1}}$ is negative , the $3_{rd}$ and $(N-2)_{th}$ sites will not support edge states. Consequently, the electron population at these sites continuously varies across the entire filling range, mirroring the behavior of bulk states. Conversely, for positive $\Delta_{J_{3}-J_{1}}$, a pair of edge states appears at the $3_{rd}$ and $(N-2)_{th}$ sites. This scenario, as illustrated in Fig. 3(b), leads to a sharp change in the electron population at half-filling. By focusing on the combined behavior of these four potential edge state sites and considering the electron filling, we observe that only all four sites support topological edge state does a sharp population change occur at half-filling, as shown in Fig. 3(c).
In the strong interaction limit (large Hubbard $U$), the Coulomb repulsion between spin-up and spin-down electrons leads to the emergence of a Mott gap in the energy spectrum. This is reflected in the addition energy spectrum, $E_{ad}(n)=E(n)-E(n-1)$ where $n$ represents the number of electrons. As expected for the Hubbard model, $E_{ad}(n)$ separates the lower and upper bands, as illustrated in Fig. 2(a) and 2(b). Notably, when Hubbard $U$ is sufficiently large, the edge states begin to be filled with electrons at quarter-filling. When considering different electron filling at this time, there will be a difference from no interaction, that is, electrons begin to occupy the edge state at quarter-filling as shown in Fig. 3(d) $\sim$ 3(f). This discrepancy arises due to the inherent gapless nature of symmetry-protected topological edge states. When the system is quarter-filling with large Hubbard $U$, introducing an additional electron at any bulk site incurs a significant energy penalty due to the Hubbard $U$ interaction. And this additional electron can occupy the edge sites without costing energy.
There are some wrinkles on the surface in Fig. 3(d) $\sim$ 3(f). This is due to the fact that $3_{rd}$ and $(N-2)_{th}$ sites to support new topological edge states following a change in the topological number $\mathcal{W}$, which allows for an additional electron to occupy these two sites without costing additional energy. And filling the sites of $1_{st}$ and $N_{th}$ extends backward by one electron. As established in Fig. 1(b), the system undergoes a topological phase transition at $J_{3}=J_{1}$, corresponding to $\Delta_{J_{3}-J_{1}}=0$. However, the wrinkles on the surface in Fig. 3(d) $\sim$ 3(f) deviate from zero. This is actually the result of the finite-size effects in the system.
@@ III.2 Bulk-edge correspondence
Figure 4: CP-AFQMC results of next-nearest neighbour hopping SSHH model with OBC at the half-filling with $J_{11}=J_{33}=0$. (a) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{3}-J_{1}}=0.75$, topological edge states at [$1_{st},3_{rd},(N-2)_{th},N_{th}$] sites. (b) Same as (a) with $\Delta_{J_{3}-J_{1}}=0$ and there is a long-range anti-ferromagnetic order. (c) Same as (a) with $\Delta_{J_{3}-J_{1}}=-0.75$, and topological edge states at [$1_{st},N_{th}$] sites. (d) The entanglement Rényi entropy between subsystem [$3_{rd},(N-2)_{th}$] sites and the remaining sites; (e) Same as (d) but subsystem [$1_{st},N_{th}$] sites and (f) with subsystem [$1_{st},3_{rd},(N-2)_{th},N_{th}$] sites. In (d) $\sim$ (f), the position where the transition occurs is closer to $\Delta_{J_{3}-J_{1}}=0$ than ED in Appendix C, which means that the finite-size effects have less impact of the conclusion.
@@
We choose $U=10,\ J_{11}=J_{33}=0$ for the investigation of the impact of $\Delta_{J_{3}-J_{1}}$ on the system , as illustrated in Fig. 4. In this configuration, $1_{st}$ and $N_{th}$ sites consistently support topological edge states. Additionally, for $\Delta_{J_{3}-J_{1}}>0$, $3_{rd}$ and $(N-2)_{th}$ sites will also support topological non-trivial edge states. These edge states are decoupled from the bulk, and each pair of edge states existing in either a singlet state |$\ket{S}=(\ket{\uparrow\downarrow}-\ket{\downarrow\uparrow})/\sqrt{2}$ or one of the three triplet states |$\ket{T_{0}}=(\ket{\uparrow\downarrow}+\ket{\downarrow\uparrow})/\sqrt{2}$, |$\ket{T_{1}}=\ket{\uparrow\uparrow}$ and |$\ket{T_{2}}=\ket{\downarrow\downarrow}$. All there states are zero-energy and gapless. When $\Delta_{J_{3}-J_{1}}<0$, the ground state of half-filling exhibits fourfold degeneracy, and for $\Delta_{J_{3}-J_{1}}>0$, the degeneracy of the ground state at half-filling increases to $16$, as confirmed by the exact diagonalization results presented in Appendix C. The existence of Hubbard $U$ reduces the ground state degeneracy. This occurs because the Hubbard $U$ prohibits the occupancy of a single site by both a spin-up and a spin-down electron.
@@
Due to the $1_{st}$ and $N_{th}$ sites can always support topological edge states, the $1_{st}$ site only has spin correlation with $N_{th}$ site as shown in Fig. 4(a) $\sim$ 4(c). When $\Delta_{J_{3}-J_{1}}>0$ in Fig. 4(a), a similar long-range spin correlation emerges between $3_{rd}$ site and $(N-2)_{th}$ site same as $1_{st}$ site and $N_{th}$ site. This is the consequence of $3_{rd}$ and $(N-2)_{th}$ which can also support topological non-trivial edge states. Conversely, when $3_{rd}$ and $(N-2)_{th}$ sites only support topological trivial states, long-range spin correlation are no longer observed, and a long range anti-ferromagnetic order is shown in Fig. 4(b). This observation is consistent with Hubbard model’s conclusion in the large $U$ limit, which can be attributed to second-order perturbation theory.
Long-range correlation effect can be observed not only through spin correlation, but also by quantifying entanglement entropy. The paired appearance of topological edge states itself means the existence of near-zero entanglement. Since the positions of edge states are located at the two boundaries of the system, the formation of topological edge states can be determined by calculating the long-range entanglement entropy between the edge sites and bulk sites.
@@
The presence of topological edge states can be judged using the second-order Rényi entanglement entropy. When topological edge states exist, the entanglement between the sites where the topological edge states appear and the remaining system will be significantly reduced, as illustrated in Fig. 4(d) $\sim$  4(f). Since $1_{st},N_{th}$ sites can always support topological edge states, leading to consistently near-zero entanglement with the bulk, as depicted in Fig. 4(e). Conversely, when $\Delta_{J_{3}-J_{1}}$ is changed until $3_{rd},(N-2)_{th}$ can support a pair of topological edge states, the entanglement entropy will also decrease to close to 0, otherwise the entanglement is larger than 0. Notably, the observed decrease in entanglement entropy does not occur precisely at $\Delta_{J_{3}-J_{1}}=0$ due to finite-size effects within the system. This effect will decrease as the system size increases. This observation aligns with the exact diagonalization results presented in the Appendix C. As larger system sizes calculated by CP-AFQMC, the position where the entanglement entropy decreases is closer to $\Delta_{J_{3}-J_{1}}=0$.
@@ III.3 Magnetic-field-induced transition at quarter-filling
Figure 6: Exact diagonalization results of next-nearest neighbour hopping SSHH model with OBC at the quarter-filling with $J_{11}=J_{33}=0$ or $J_{1}=J_{3}=0$ for (b)(d)(f). (a) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{3}-J_{1}}=-0.75,\ J_{11}=J_{33}=0$, topological edge states are [$1_{st},N_{th}$] sites, and $\omega$ is degeneracy of ground state. (b) Same as (a) with $\Delta_{J_{3}-J_{1}}=0.75$ and topological edge states are [$1_{st},3_{rd},(N-2)_{th},N_{th}$] sites. (c) Rényi entanglement entropy between the edges [$3_{rd},(N-2)_{th}$] sites and the bulk for various field strength and $\Delta_{J_{3}-J_{1}}$. And each contour can separate with different $S_{z}$, where $S_{z}=2$ means the maximally ferromagnetic ground state. (d) Same as (c) but with topological [$2_{nd},(N-1)_{th}$] sites. (e) Critical field strength at various Hubbard $U$ and $\Delta_{J_{3}-J_{1}}$. (f) Same as (e) but changing with $\Delta_{J_{33}-J_{11}}$. If we do a $1D$ slice for $\Delta_{J_{3}-J_{1}}=0.75$ in (e) or $\Delta_{J_{33}-J_{11}}=0.75$ in (f), there will be a decrease in the critical field with increasing Hubbard $U$. (a) and (b) for a $1D$ chain with $N=12$, and (c) $\sim$ (f) for a 1D chain with $N=8$.
Figure 7: CP-AFQMC results of next-nearest neighbour hopping SSHH model with OBC at the quarter-filling with $J_{11}=J_{33}=0$, and $U=10$ for (a) $\sim$ (e). (a) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{3}-J_{1}}=-0.75$, topological edge states at [$1_{st},N_{th}$] sites. (b) Same as (a) with $\Delta_{J_{3}-J_{1}}=0.75$ and topological edge states at [$1_{st},3_{rd},(N-2)_{th},N_{th}$] sites. (c) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{3}-J_{1}}=-0.75$ above the critical magnetic field. (d) Same as (c) but $\Delta_{J_{3}-J_{1}}=0.75$. (e) Rényi entanglement entropy with changing external magnetic field and hopping amplitude difference $\Delta_{J_{3}-J_{1}}$. Each point has error bar in both x and y direction, and the dotted line is a schematic line connecting different $\Delta_{J_{3}-J_{1}}$. (f) Critical field strength at various of the on-site interaction and hopping amplitude difference and it can be seen that for a certain $\Delta_{J_{3}-J_{1}}$, the value of the critical magnetic field will decrease as $U$ increases.
@@
The Hamiltonian describing the system under a magnetic field is given by: $H_{1}=H-(E_{B}/2)\Sigma_{i}(\hat{n}_{i\uparrow}^{s}-\hat{n}_{i\downarrow}^{s})$ where $E_{B}=g\mu_{B}B$, $\mu_{B}$ is the Bohr magneton and $g$ is the g-factor of the electron in the material. According to the above discussion, the system exhibits no long-range spin correlation at quarter-filling. However, a sufficiently strong magnetic field can induce a transition to a maximally ferromagnetic ground state, as evidenced by the emergence of long-range spin correlation in Fig. 6(c) and 6(d). This can be attributed to the magnetic field effectively transforming the quarter-filling state into a quasi half-filling state at high field strengths, which is similar to the true half-filling under zero magnetic field. Additionally, the Rényi entropy drops to 0 for $\Delta_{J}>0$ due to the formation of the new topological edge states.
@@
To achieve a maximally ferromagnetic state, the last spin-down electron needs to be pumped to the next unoccupied spin-up single-particle state. The existence of Hubbard $U$ causes that at quarter-filling, the energy of only one spin-down electron is higher than $U=0$ case. Consequently, a smaller critical magnetic field is required to overcome the energy barrier for flipping this spin-down electron to a spin-up state. This will make the critical magnetic field that needs to be applied in the experiment ten times smaller, as demonstrated in Fig. 6(e) and 6(f).
@@
Although there will be a sign problem at quarter-filling, we can calculate the above results under different electron numbers using CP-AFQMC for a larger $1D$ chain with $N=28$, as shown in Fig. 7. The presence of a magnetic field introduces topological edge states component to the ground state at quarter-filling. However, the trial wave function at quarter-filling obtained from a single-particle model does not contain edge component. Therefore, we can no longer simply use the single-particle wave function as the trial wave function. Here, we propose a new trial wave function by linearly superimposing the single-particle wave function at quarter-filling with the topological edge states, and this new trial wave function at quarter-filling can be used to reduce the system’s bias.
In Fig. 7(a) $\sim$ 7(d), we employ the formula (11) to calculate the spin correlation, we can see the conclusion consistent with Fig. 6, the long-range AFM order manifested under the critical magnetic field and the long-range spin correlation above the critical magnetic field. Subsequently, the Rényi entanglement entropy of the system is calculated using formulas (10) as depicted in Fig. 7(e). As can be observed, for $\Delta_{J_{3}-J_{1}}<0$, the entanglement entropy remains nonzero regardless of the applied magnetic field. Conversely, when $\Delta_{J_{3}-J_{1}}>0$, the formation of long-range spin correlation leads to a drop in entanglement entropy towards zero as the magnetic field strength increases. This behavior is consistent with Fig. 6(c) and 6(d), the entanglement entropy within each contour is a constant value. The calculated entanglement entropy for different spin configurations is connected by a polyline fit. And the critical magnetic field, corresponding to the formation of the maximally ferromagnetic state, is defined as the magnetic field strength at which the entanglement entropy reaches its minimum for different values of $\Delta_{J_{3}-J_{1}}$. As illustrated in Fig. 7(f), the relationship between the critical magnetic field and the Hubbard $U$ is determined using CP-AFQMC. Similarly, a substantial decrease in the critical field with increasing Hubbard $U$ is observed. The case $J_{1}=J_{3}=0$ is detailed in the Appendix D.

### v2

@@ I Introduction
However, in many experiments setups, the interaction between electrons dominates over the hopping amplitude [23, 24, 25], rendering the independent-electron approximation invalid. Topological descriptions of strongly correlated systems are an emerging area of research with significant potential [26, 27, 28]. In interacting systems, the bulk-edge correspondence is modified as the relation between bulk topological property and the many-body, ground-state degeneracy on the edge [29]. Researchers have also developed a range of metrics to evaluate characterize topological edge states, including the Green function method [30, 31, 32], entanglement entropy between topological edges and bulk states [33, 34], and topological phase transition induced by magnetic flux changes [35].
Su-Schrieffer-Heeger-Hubbard (SSHH) model with the nearest neighbour hopping has been extensively studied. However, the one with next-nearest neighbour hopping remains largely unexplored. Here, we discuss the ground-state properties of a one-dimensional SSHH model with next-nearest neighbour hopping. In the absence of Hubbard interaction, this model can be well described by the single-particle theory and exhibits four distinct topological phases ($\mathcal{W}=-1,0,1,2$) without breaking chiral symmetry [36, 37]. According to the bulk-edge correspondence, these topological phases are characterized by the degeneracy of zero-energy edge modes. However, in the presence of onsite Hubbard interaction, we should consider the relation between bulk topological property and the many-body ground-state. We evaluate long-range spin correlation and the second-order Rényi entropy of the ground state, and determine the topological state of the system by measuring the correlation between the boundary and bulk. The presence of Coulomb interaction leads to a Mott gap at half-filling and dimerization at quarter-filling [38]. To mitigate the influence of finite site effects on our results, we employ the constrained-path auxiliary-field quantum Monte Carlo (CP-AFQMC) algorithm. Furthermore, the applicability of constrained-path method extends beyond the half-filling case [34] where there is no sign problem. Our analysis reveals that the system exhibits long-range spin correlation and near-zero Rényi entropy at half-filling, indicative of a topological order state. Furthermore, a long-range antiferromagnetic order emerges at quarter-filling. The application of an external magnetic field disrupts this antiferromagnetic order, leading to a restoration of long-range spin correlation and near-zero Rényi entropy.
@@ III Result
Figure 3: Population number of the edge states for $N=8$ sites chain with $J_{11}=J_{33}=0$, and $n$ is the number of electrons. (a)-(c) for $U=0$ and (d)-(f) for $U=10$. According to Fig. 1(e), the edge states appear at the first boundary ($1_{st}$ and $N_{th}$) sites or the third-boundary ($3_{rd}$ and $(N-2)_{th}$) sites. (a), (d) $n_{edge1}$ is the sum of electron number on the $1_{st}$ and $N_{th}$ sites; (b), (e) $n_{edge2}$ is the sum of electron number on the $3_{rd}$ and $(N-2)_{th}$ sites; (c), (f) $n_{edge3}$ is the sum of electron number on above four sites. In the non-trivial phases without Hubbard $U$, the edge population increases sharply at half-filling, while with Hubbard $U$ the edge population increases sharply at quarter-filling and there-quarter-filling. And in the trivial phase, the edge population increases gradually. There are some wrinkles on the surface in (d)-(f), which is due to the finite-size effects.
Figure 4: CP-AFQMC results of next-nearest neighbour hopping SSHH model with OBC at the half-filling with $J_{11}=J_{33}=0$. (a) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{3}-J_{1}}=0.75$, topological edge states at [$1_{st},3_{rd},(N-2)_{th},N_{th}$] sites. (b) Same as (a) with $\Delta_{J_{3}-J_{1}}=0$ and there is a long-range anti-ferromagnetic order. (c) Same as (a) with $\Delta_{J_{3}-J_{1}}=-0.75$, and topological edge states at [$1_{st},N_{th}$] sites. (d) The entanglement Rényi entropy between subsystem [$3_{rd},(N-2)_{th}$] sites and the remaining sites; (e) Same as (d) but subsystem [$1_{st},N_{th}$] sites and (f) with subsystem [$1_{st},3_{rd},(N-2)_{th},N_{th}$] sites. In (d)-(f), the position where the transition occurs is closer to $\Delta_{J_{3}-J_{1}}=0$ than ED in Appendix C, which means that the finite-size effects have less impact of the conclusion.
Figure 6: Exact diagonalization results of next-nearest neighbour hopping SSHH model with OBC at the quarter-filling with $J_{11}=J_{33}=0$ or $J_{1}=J_{3}=0$ for (b)(d)(f). (a) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{3}-J_{1}}=-0.75,\ J_{11}=J_{33}=0$, topological edge states are [$1_{st},N_{th}$] sites, and $\omega$ is degeneracy of ground state. (b) Same as (a) with $\Delta_{J_{3}-J_{1}}=0.75$ and topological edge states are [$1_{st},3_{rd},(N-2)_{th},N_{th}$] sites. (c) Rényi entanglement entropy between the edges [$3_{rd},(N-2)_{th}$] sites and the bulk for various field strength and $\Delta_{J_{3}-J_{1}}$. And each contour can separate with different $S_{z}$, where $S_{z}=2$ means the maximally ferromagnetic ground state. (d) Same as (c) but with topological [$2_{nd},(N-1)_{th}$] sites. (e) Critical field strength at various Hubbard $U$ and $\Delta_{J_{3}-J_{1}}$. (f) Same as (e) but changing with $\Delta_{J_{33}-J_{11}}$. If we do a $1D$ slice for $\Delta_{J_{3}-J_{1}}=0.75$ in (e) or $\Delta_{J_{33}-J_{11}}=0.75$ in (f), there will be a decrease in the critical field with increasing Hubbard $U$. (a) and (b) for a $1D$ chain with $N=12$, and (c)-(f) for a 1D chain with $N=8$.
Figure 7: CP-AFQMC results of next-nearest neighbour hopping SSHH model with OBC at the quarter-filling with $J_{11}=J_{33}=0$, and $U=10$ for (a)-(e). (a) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{3}-J_{1}}=-0.75$, topological edge states at [$1_{st},N_{th}$] sites. (b) Same as (a) with $\Delta_{J_{3}-J_{1}}=0.75$ and topological edge states at [$1_{st},3_{rd},(N-2)_{th},N_{th}$] sites. (c) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{3}-J_{1}}=-0.75$ above the critical magnetic field. (d) Same as (c) but $\Delta_{J_{3}-J_{1}}=0.75$. (e) Rényi entanglement entropy with changing external magnetic field and hopping amplitude difference $\Delta_{J_{3}-J_{1}}$. Each point has error bar in both x and y direction, and the dotted line is a schematic line connecting different $\Delta_{J_{3}-J_{1}}$. (f) Critical field strength at various of the on-site interaction and hopping amplitude difference and it can be seen that for a certain $\Delta_{J_{3}-J_{1}}$, the value of the critical magnetic field will decrease as $U$ increases.
@@ III.1 Edge population
The difference between nearest neighbour hopping and next-nearest hopping $\Delta_{J_{3}-J_{1}}$ affects the topological properties of $3_{rd}$ and $(N-2)_{th}$ sites according to Fig. 1(c). When $\Delta_{J_{3}-J_{1}}$ is negative , the $3_{rd}$ and $(N-2)_{th}$ sites do not support edge states. Consequently, the electron population at these sites continuously varies across the entire filling range, mirroring the behavior of bulk states. Conversely, for positive $\Delta_{J_{3}-J_{1}}$, a pair of edge states appears at the $3_{rd}$ and $(N-2)_{th}$ sites. This scenario, as illustrated in Fig. 3(b), leads to a sharp change in the electron population at half-filling. By focusing on the combined behavior of these four potential edge state sites and considering the electron filling, we observe that only all four sites support topological edge state does a sharp population change occur at half-filling, as shown in Fig. 3(c).
In the strong interaction limit (large Hubbard $U$), the Coulomb repulsion between spin-up and spin-down electrons leads to the emergence of a Mott gap in the energy spectrum. This is reflected in the addition energy spectrum, $E_{ad}(n)=E(n)-E(n-1)$ where $n$ represents the number of electrons. As expected for the Hubbard model, $E_{ad}(n)$ separates the lower and upper bands, as illustrated in Fig. 2(a) and 2(b). Notably, when Hubbard $U$ is sufficiently large, the edge states begin to be filled with electrons at quarter-filling. When considering different electron filling at this time, there will be a difference from no interaction, that is, electrons begin to occupy the edge state at quarter-filling as shown in Fig. 3(d)-3(f). This discrepancy arises due to the inherent gapless nature of symmetry-protected topological edge states. When the system is quarter-filling with large Hubbard $U$, introducing an additional electron at any bulk site incurs a significant energy penalty due to the Hubbard $U$ interaction. This additional electron can occupy the edge sites without costing energy.
There are some wrinkles on the surface in Fig. 3(d)-3(f). This is due to the fact that $3_{rd}$ and $(N-2)_{th}$ sites to support new topological edge states following a change in the topological number $\mathcal{W}$, which allows for an additional electron to occupy these two sites without costing additional energy. And filling the sites of $1_{st}$ and $N_{th}$ extends backward by one electron. As established in Fig. 1(b), the system undergoes a topological phase transition at $J_{3}=J_{1}$, corresponding to $\Delta_{J_{3}-J_{1}}=0$. However, the wrinkles on the surface in Fig. 3(d)-3(f) deviate from zero. This is actually the result of the finite-size effects in the system.
@@ III.2 Bulk-edge correspondence
We choose $U=10,\ J_{11}=J_{33}=0$ for the investigation of the impact of $\Delta_{J_{3}-J_{1}}$ on the system , as illustrated in Fig. 4. In this configuration, $1_{st}$ and $N_{th}$ sites consistently support topological edge states. Additionally, for $\Delta_{J_{3}-J_{1}}>0$, $3_{rd}$ and $(N-2)_{th}$ sites also support topological non-trivial edge states. These edge states are decoupled from the bulk, and each pair of edge states existing in either a singlet state |$\ket{S}=(\ket{\uparrow\downarrow}-\ket{\downarrow\uparrow})/\sqrt{2}$ or one of the three triplet states |$\ket{T_{0}}=(\ket{\uparrow\downarrow}+\ket{\downarrow\uparrow})/\sqrt{2}$, |$\ket{T_{1}}=\ket{\uparrow\uparrow}$ and |$\ket{T_{2}}=\ket{\downarrow\downarrow}$. All there states are zero-energy and gapless. When $\Delta_{J_{3}-J_{1}}<0$, the ground state of half-filling exhibits fourfold degeneracy, and for $\Delta_{J_{3}-J_{1}}>0$, the degeneracy of the ground state at half-filling increases to $16$, as confirmed by the exact diagonalization results presented in Appendix C. The existence of Hubbard $U$ reduces the ground state degeneracy. This arises due to the exclusion principle enforced by U, forbidding the co-occupation of a single site by electrons with opposing spins.
@@
As the $1_{st}$ and $N_{th}$ sites can always support topological edge states, the $1_{st}$ site only has spin correlation with $N_{th}$ site as shown in Fig. 4(a)-4(c). When $\Delta_{J_{3}-J_{1}}>0$ [Fig. 4(a)], a similar long-range spin correlation emerges between $3_{rd}$ site and $(N-2)_{th}$ site same as $1_{st}$ site and $N_{th}$ site. This is because the $3_{rd}$ and $(N-2)_{th}$ can also support topological non-trivial edge states. Conversely, when $3_{rd}$ and $(N-2)_{th}$ sites only support topological trivial states, long-range spin correlation are no longer observed, and a long range anti-ferromagnetic order is shown in Fig. 4(b). This observation is consistent with Hubbard model’s conclusion in the large $U$ limit, which can be attributed to second-order perturbation theory.
Long-range correlation effect can be observed not only through spin correlation, but also by quantifying entanglement entropy. The paired appearance of topological edge states means the existence of near-zero entanglement. Since the positions of edge states are located at the two boundaries of the system, the formation of topological edge states can be determined by calculating the long-range entanglement entropy between the edge sites and bulk sites.
@@
The presence of topological edge states can be judged using the second-order Rényi entanglement entropy. When topological edge states exist, the entanglement between the sites where the topological edge states appear and the remaining system will be significantly reduced, as illustrated in Fig. 4(d)- 4(f). Since $1_{st},N_{th}$ sites can always support topological edge states, leading to near-zero entanglement with the bulk, as depicted in Fig. 4(e). Conversely, when $\Delta_{J_{3}-J_{1}}$ is changed until $3_{rd},(N-2)_{th}$ can support a pair of topological edge states, the entanglement entropy also decrease close to zero. In all other configurations, the entanglement remains non-zero. Notably, the observed decrease in entanglement entropy does not occur precisely at $\Delta_{J_{3}-J_{1}}=0$ due to finite-size effects within the system. This effect will decrease as the system size increases. This observation aligns with the exact diagonalization results presented in the Appendix C. As larger system sizes calculated by CP-AFQMC, the position where the entanglement entropy decreases is closer to $\Delta_{J_{3}-J_{1}}=0$.
@@ III.3 Magnetic-field-induced transition at quarter-filling
The Hamiltonian describing the system under a magnetic field is given by $H_{1}=H-(E_{B}/2)\Sigma_{i}(\hat{n}_{i\uparrow}^{s}-\hat{n}_{i\downarrow}^{s})$ with $E_{B}=g\mu_{B}B$, where $\mu_{B}$ is the Bohr magneton and $g$ is the g-factor of the electron in the material. According to the above discussion, the system exhibits no long-range spin correlation at quarter-filling. However, a sufficiently strong magnetic field can induce a transition to a maximally ferromagnetic ground state, as evidenced by the emergence of long-range spin correlation in Fig. 6(c) and 6(d). This can be attributed to the magnetic field effectively transforming the quarter-filling state into a quasi half-filling state at high field strengths, which is similar to the true half-filling under zero magnetic field. Additionally, the Rényi entropy drops to zero for $\Delta_{J}>0$ due to the formation of new topological edge states.
@@
To achieve a maximally ferromagnetic state, the last spin-down electron needs to be pumped to the next unoccupied spin-up single-particle state. The existence of Hubbard $U$ causes that at quarter-filling, the energy of only one spin-down electron is higher than $U=0$ case. Consequently, a smaller critical magnetic field is required to overcome the energy barrier for flipping this spin-down electron to a spin-up state. Since it is very difficult to realize a strong magnetic field in experiment, the existence of Hubbard $U$ allows us to observe the above transformation under a much smaller magnetic field, as demonstrated in Fig. 6(e) and 6(f).
@@
Although there will be a sign problem at quarter-filling, we can calculate the above results under different electron numbers using CP-AFQMC for a larger $1D$ chain with $N=28$, as shown in Fig. 7. The presence of a magnetic field introduces topological edge states component to the ground state at quarter-filling. However, the trial wave function at quarter-filling obtained from a single-particle model does not contain edge component. Therefore, we can no longer simply use the single-particle wave function as the trial wave function. Here, we propose a new trial wave function by linearly superimposing the single-particle wave function at quarter-filling with the topological edge states. This new trial wave function at quarter-filling can be used to reduce the system’s bias.
In Fig. 7(a)-7(d), we employ the formula (11) to calculate the spin correlation. The results consistent with Fig. 6, that is, the long-range AFM order manifested under the critical magnetic field and the long-range spin correlation above the critical magnetic field. Subsequently, the Rényi entanglement entropy of the system is calculated using formulas (10), as depicted in Fig. 7(e). We can observe that for $\Delta_{J_{3}-J_{1}}<0$, the entanglement entropy remains nonzero regardless of the applied magnetic field . Conversely, when $\Delta_{J_{3}-J_{1}}>0$, the formation of long-range spin correlation leads to a drop in entanglement entropy towards zero as the magnetic field strength increases. This behavior is consistent with Fig. 6(c) and 6(d), the entanglement entropy within each contour is a constant value. The calculated entanglement entropy for different spin configurations is connected by a polyline fit. The critical magnetic field, corresponding to the formation of the maximally ferromagnetic state, is defined as the magnetic field strength at which the entanglement entropy reaches its minimum for different values of $\Delta_{J_{3}-J_{1}}$. As illustrated in Fig. 7(f), the relationship between the critical magnetic field and the Hubbard $U$ is determined using CP-AFQMC. Similarly, a substantial decrease in the critical field with increasing Hubbard $U$ is observed. The case $J_{1}=J_{3}=0$ is detailed in the Appendix D.